The number of credits being taken by a sample of 13​ full-time college students are listed below. Find the​ mean, median, and mode of the​ data, if possible. If any measure cannot be found or does not represent the center of the​ data, explain why.

88  

1010  

1111  

1111  

88  

99  

77  

77  

77  

77  

77  

77  

88  
Answer #1

Given data is 8,10,11,11,8,9,7,7,7,7,7,7,8

arranging data in ascending order :_ 7,7,7,7,7,7,8,8,8,9,10,11,11

number of data items is 13, i.e. odd value

Median = (n+1)/2 term, where n = 13

median = (13+1)/2 = 7th term, which is 8

and mode = 7 (because 7 has highest frequency of 6)